OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (–9, 6)?

OpenStudy (anonymous):

\[y = -\frac{ x }{ 3 } + 3\]

OpenStudy (waleed_imtiaz):

Just use two point form.... Can U do this ?

OpenStudy (waleed_imtiaz):

y-y1/y2-y1 = x-x1/x2-x1..... Apply this..... U will get the equation .......